Sinopec Group 4312 (a) Give the f
(b) Which companies are outliers according to the 1.5 x IOR rule? (Select all that apply) E Ford Motor Wal-Mart Stores Allianz Exxon Mobil e Royal Dutch Shell Sinopec Group Hon Hai Precision Industry Industrial & Commercial Bank of China
Answer #1

sort the data from lowest to highest

a)

minimum=   130.8
quartile , Q1 = 0.25(n+1)th value=   8 th value =   146.9  

Median=0.5(n+1)th value =    16   th value=   169.6

Quartile , Q3 = 0.75(n+1)th value=   24th value=   268.7

maximum =    484.7
b)

IQR = Q3-Q1 =    121.8  
      
1.5IQR =    182.7  
      
lower bound=Q1-1.5IQR=   -35.8  
      
upper bound=Q3+1.5IQR=   451.4  
      
outlier =values outside lower bound and upper bound      
so, there are three outliers

ROYAL DUTCH

WAL MART STORES

EXXON MOBIL